OpenStudy (anonymous):

find the critical point(s) of f(x, y) = x^2 + y^2 + 2xy and classify them as minimum, maximum or saddle points

OpenStudy (anonymous):

i'll get a start on it: Fx = 2x + 2y Fy = 2x + 2y so Fx = 0 when y = -x so critical points lie on this line

OpenStudy (anonymous):

anyone able to help? i get stuck on the Hessian

OpenStudy (anonymous):

To find the critical points, find the points at which \(f_x=0 \text{ and } f_y=0\). In our case, we have \(f_x=2x+2y\text{ and } f_y=2y+2x\). Therefore, the critical points are all the points on the line \(y=-x\).

OpenStudy (anonymous):

yep, i got that far. i can't figure out if it's a maximum or minimum

OpenStudy (anonymous):

It can't be maximum; f(x,y) can go to infinity.

OpenStudy (anonymous):

yes, but to find out whether it's a maximum or minimum, you find the hessian matrix

OpenStudy (anonymous):

This test will fail here because \(f_{xx}f{yy}-[f{xy}]^2=0\). You can see it that it's minimum by writing \(f(x,y)=(x+y)^2\). That shows that f is at least \(0\), and that happens at all points where \(y=-x\).

OpenStudy (anonymous):

see that*

OpenStudy (anonymous):

yes, i can see it's a minimum, but how do i prove it?

OpenStudy (anonymous):

shouldn't the test be succeeding every time?

OpenStudy (anonymous):

No, the test is inconclusive when \(f_{xx}f_{yy}-f^2_{xy}=0\).

OpenStudy (anonymous):

so what do you have to do if it's inconclusive?

OpenStudy (anonymous):

I'm not sure if there's another test you can use here.

OpenStudy (anonymous):

ok, what about f(x,y) = e^(1+x^2-y^2)? answer says there's a saddle point at (0,0)

OpenStudy (anonymous):

Fx = 2x*e^(1+x^2 - y^2) Fy = -2y*e^(1+x^2 - y^2) so Fx = 0 when x = 0 and Fy = 0 when y = 0 so critical point at (0,0)

OpenStudy (anonymous):

Fxy = -4xy*e^(1+x^2 - y^2) so Fxy(0,0) = 0 Fxx = 4x^2*e^(1+x^2-y^2) so Fxx(0,0) = 0 Fyx = -4xy*e^(1+x^2 - y^2) so Fyx(0,0) = 0 Fyy(0,0) = 0 so determinant is once again 0, so the test is inconclusive

OpenStudy (anonymous):

so how the hell do we know it's a saddle point?

OpenStudy (anonymous):

This would probably help http://mathworld.wolfram.com/ExtremumTest.html

OpenStudy (anonymous):

that's in one variable though

OpenStudy (anonymous):

oh sorry!

OpenStudy (anonymous):

thanks for your help anyway :)